I have a bag of N white marbles. I paint 20 of the marbles black. Later, my sister pulls out 30 marbles, and I tell her that my best guess is that 12 of them will be black. How many marbles are in the bag
Answer #1

Out of the 30 balls which are pulled out, it can be almost surely said that 12 of them are black, which means that the remaining 18 balls are white. Therefore, black and white balls are in the ratio 12:18 = 2:3

Now, if we know that we have 20 black balls then

20/x =2/3, where X = No. of white balls.

Therefore, x = 20×3/2 = 30 white balls..

Total marbles in the bag = 20 black + 30 white = 50.

> Why the ratio is 20/x not 20/(x-20)? Because you did 12/(30-12) = 12/18, then why are not you doing the same for 20/x?

MillionTrade Tue, Feb 8, 2022 10:54 PM